OpenStudy (cometailcane):

How many H atoms are in 2.50g of (NH4)2SO4?

OpenStudy (cometailcane):

So I found that the molar mass of \[(NH _{4})_{2}SO _{4}\] is 132.154 g/mol

OpenStudy (cometailcane):

I set up this equation: \[2.50 g (\frac{ 1 mol }{ 132.154 g }) = 0.018917324\]

OpenStudy (cometailcane):

Then I did \[1.89 \times 10^{-2} mol of H (\frac{ 8.064 g }{ mol of H }) = 1.52 \times 10^{-1} g\]

OpenStudy (cometailcane):

Is what I am doing correct? And how do I convert my answer into atoms?

OpenStudy (sapphiremoon):

Ah. Found an error. What you have after your first equation is how many moles of ammonium? sulfate you have (just finished nomenclature, not an expert yet lol). After that you'll want to do:\[mol (NH_{4})_2SO_{4} \times \frac{ 8 mol H }{ mol (NH_{4})_2SO_{4}}\] which will give you the moles of hydrogen. Then you don't need to convert to grams again, just multiply the mol H by Avogadro's number to get the number of atoms. Any more questions?

OpenStudy (cometailcane):

Are you saying do \[1.89 \times 10^{-2} \times 8 mol H\] and then multiply by Avogadro's number?

OpenStudy (sapphiremoon):

Yep, that's what it comes out to, I believe.